For the shaded and shown, determine the location of the cartold - R 105 of ROS 200mm Dimensions in
Answer #1

2105 210 - - - - - - 2uoma - Slino Share l kom I (mn) Rectangle . Arca (mm) | A1: :210x240 | x1 = 24/ 350400 1 = 120 14,=210- - -- cancroid: ī - AUXH A2X2-A3X3 AHA2-A3 = 50400 (120) +17318.029(284.563) -663-614 X(27.5363 50400 +17318.029 - 6636.614